# Point-In-Context: Understanding Point Cloud via In-Context Learning

### v1

@@ 3.1 Modeling In-Context Learning in 3D Point Cloud
In-Context Learning in 2D. Previous works, such as Visual Prompt [2] and Painter [13] combine two pairs of images that perform the same task into a grid-like image and randomly mask portions, following MAE [26]. During training, these models [2, 13] take two pairs of images $I_{i}$ and $I_{j}$ as inputs and form two pairs, including a reference pair (also termed prompt), $R^{k}_{i}=\{I_{i},T^{k}_{i}\}$ and a query pair, $Q^{k}_{j}=\{I_{j},T^{k}_{j}\}$. $R^{k}_{i}$ is the task prompt containing one image $I_{i}$ and its corresponding target $T^{k}_{i}$. $Q^{k}_{j}$ represent the current input image $I_{j}$ and its target $T^{k}_{j}$, which perform the same task with the prompt pair. Here, $k$ represents the task index. When $i=j$, we term the prompt as an ideal prompt. During inference, only task example pairs and a query image are provided. They are combined into a grid-like image with a quarter mask to mask $T^{k}_{j}$, and a pre-trained model is used to restore the missing parts, as shown in Fig. 1 (b).
Figure 2: (a) The pre-training pipeline used in previous works. When performing Masked Point Modeling (MPM), these works [67, 27, 29] use the center position of the target patches for position embedding, which results in information leakage. (b) Difference between 2D images and 3D point clouds. For 2D images, the semantic information at corresponding pixel positions in the input and target images is consistent. However, after performing grouping operations on 3D point clouds, the order of points will inevitably be disrupted, resulting in inconsistency between the point sequences of the input and output point clouds. (c) Joint Sampling module involves recording the indices of sampled center points and employing the K-nearest neighbor strategy to sample both the input and target point clouds concurrently.
Modeling In-Context Learning in 3D. Following the same spirit of 2D in-context learning, for the first time, we design a similar procedure for 3D in-context learning. During training, each input sample contains two pairs of point clouds, including prompt point clouds $R^{k}_{i}=\{P_{i},T^{k}_{i}\}$ and query point clouds $Q^{k}_{j}=\{P_{j},T^{k}_{j}\}$, which perform the same task as in 2D in-context learning. Each pair consists of an input point cloud and its corresponding target point cloud for the given task. Similar to PointMAE [27], we adopt the farthest point sampling (FPS) and K-nearest neighbor (KNN) techniques to convert the point clouds into a sentence-like data format. These point patches are subsequently encoded into tokens. During inference, the input point cloud is a combination of prompt and query point cloud, while the target point cloud consists of an example target along with masked tokens, as shown in Fig. 1(c) and Fig. 1(d). Consequently, we obtain predictions corresponding to the position of the mask. Given an input, the models can output corresponding targets following various prompts.
@@ 3.2 ShapeNet In-Context Datasets
To establish the first benchmark for 3D in-context learning, we carefully curate datasets and define task specifications, as there is no prior benchmark. Initially, we collect samples from representative datasets like ShapeNet [30] and ShapeNetPart [3], transforming them into the “input-target” format as stated in Sec. 3.1. To augment sample size for part segmentation, we apply random operations such as point cloud perturbation, rotation, and scaling to ShapeNetPart. This yields an extensive dataset with 217,454 samples across four tasks. Each sample includes an input point cloud and its corresponding task-specific target. We standardize the XYZ coordinates of inputs and outputs to the range between $[-1,1]$, and unify all samples as pure point clouds in $(N,3)$ shape.
@@
Registration. This task aims to restore a rotated point cloud to its initial orientation. We assume that during both training and inference, the query point cloud and the prompt are synchronized in terms of the rotation angle. To prevent interference with the denoising and reconstruction tasks, we configure the registration task’s output to include both an upright and an upside-down point cloud. We also introduce five corruption levels for this task.
@@
3.3 Human & Object Segmentation In-Context Datasets
@@
To further improve the performance and generalization of PIC to segmentation tasks, we collect more datasets related to part segmentation in point clouds. These datasets include ShapeNetPart [3], originally used for synthetic object part segmentation, Human3D [4] for human body-part segmentation, BEHAVE [5] for human-object interaction segmentation, and AKB-48 [6] for real-collected object part segmentation. These datasets contain point clouds widely used in 3D human-machine interaction scenarios [103], including human, object, and human-object interaction point clouds. We process these datasets into a format suitable for part segmentation and integrate them into large-scale Human & Object Segmentation In-Context Datasets, consisting of approximately 140k samples. Moreover, to test the model’s generalization, we exclude AKB-48 from the training set and use it as a one-shot generalization test dataset.
@@
ShapeNetPart [3] dataset consists of 16,881 shapes categorized into 16 classes, with 50 part labels. Within each class, the number of parts varies from 2 to 6. As described in Sec. 3.2, we aim to augment the challenge of the ShapeNetPart dataset and enhance its representation of real-world point cloud scenarios. To achieve this, we apply various random operations to the dataset, introducing noise to simulate imperfect point clouds commonly encountered in real-world environments. Then, we transform all part segmentation labels of a sample into points in space, each point containing XYZ coordinates. These transformed points are then aggregated to construct a point cloud with the same shape as the input point cloud. Finally, we combine the input point cloud with the labeled point cloud to create a sample in our dataset.
@@
Human3D [4] is a dataset comprising point cloud videos captured in real-world scenarios, annotated based on human body parts. Each sample of it contains two persons, both divided into 15 parts. To seamlessly integrate prompts into our proposed in-context learning framework, we divide each sample containing two individuals into two separate samples. Each sample focuses on one person with annotations for human-part segmentation. Given the complexity and difficulty in distinguishing between the original 15 body parts, we merge them into 6 categories: head, body, left arm, right arm, left leg, and right leg. This consolidation aims to enhance the semantic representation of human body parts. Additionally, we treat each frame of the dynamic video as a static sample in our dataset. This approach allows us to create a single-person human body part segmentation dataset, referred to as the enhanced version of Human3D, comprising 57,078 samples.
@@
BEHAVE [5] is an extensive dataset focusing on full-body human-object interactions, featuring multi-view RGBD frames paired with corresponding 3D SMPL and object fits. Additionally, it includes annotated contact points between humans and objects. The BEHAVE dataset encompasses roughly 15K frames capturing 8 subjects engaging in various interactions with 20 commonly encountered objects. Presented in video format, each frame of the RGBD video is transformed into point clouds, and treated as an individual sample within our dataset. These point cloud samples are annotated for segmentation, covering both the person and the object. After gathering around 15k point cloud samples, we excluded samples in which the proportion of points attributed to objects is smaller than 5% of the entire point cloud sample.
@@
AKB-48 [6] is a comprehensive-articulated object dataset containing 2,037 real-world 3D models across 48 categories. Each model is meticulously scanned from its real-life counterpart and further refined manually. Throughout the data collection, AKB-48 categorizes objects into multiple components, facilitating their utilization in part segmentation tasks. However, some categories in AKB-48 inherently pose challenges due to semantic ambiguity or confusion. For example, within the “toy” category, samples may include both dinosaurs and cars, which do not share identical labels, highlighting semantic ambiguity. Another instance is the “folding rack” category, which presents a symmetric object segmented into two or even four components despite its identical left and right sides, showcasing semantic confusion. To address these issues, we conduct filtering and preprocessing operations on the AKB-48 dataset. Firstly, categories prone to semantic confusion are removed. Furthermore, we manually consolidate components with similar semantics into a unified category. Finally, following a similar approach to the ShapeNetPart dataset, each component is labeled and merged to form a comprehensive point cloud with segmentation labels. Given AKB-48’s diverse component categories and absence of overlap with previous datasets, we consider it a one-shot segmentation generalization test dataset.
@@ 4 Methodology
In this section, we first introduce the vanilla version of PIC, named Point-In-Context-Generalist (PIC-G), a general point cloud framework capable of handling multiple tasks via the in-context learning capability. Next, to enhance the performance and generalization of PIC in segmentation tasks, we highlight the extended version of PIC, termed Point-In-Context-Segmenter (PIC-S), which is dedicated to part segmentation across multiple datasets as a part segmentation specialist.
@@
Figure 3: Overall scheme of our Point-In-Context-Generalist. Top: Training pipeline of the Masked Point Modeling (MPM) framework. During training, each sample comprises two pairs of input and target point clouds that tackle the same task. These pairs are fed into the transformer model to perform the masked point reconstruction task, which follows a random masking process. Bottom: In-context inference on multitask. Our Point-In-Context could infer results on various downstream point cloud tasks, including reconstruction, denoising, registration, and part segmentation.
4.1 Point-In-Context-Generalist Model
@@
MPM for 3D In-Context Learning. Building upon previous works [2, 13], we adopt Masked Point Modeling (MPM) for the training of point cloud understanding and introduce Point-In-Context-Generalist (PIC-G). In PIC-G, we conceptualize the local area of point clouds within the MPM as akin to image tokens within the MIM, thereby enabling us to harness the power of the transformer architecture for processing both point cloud and image data seamlessly. Illustrated in Fig. 3, our training process revolves around the objective of reconstructing the masked point patches, facilitated by the presence of input point clouds and visible target point cloud patches. During the inference stage, given a query point cloud, PIC-G will perform the specific task: reconstruction, denoising, registration, or part segmentation, according to the prompt. Subsequently, it generates target outputs corresponding to the query according to the task demonstrated by the prompt.
@@
Information Leakage. The conventional 3D MPM approach of employing position embeddings may introduce information leakage. That is because previous pre-training pipelines embed positional information with the sampled center point coordinates of all patches, even for those that are masked out (invisible), as shown in Fig. 2 (a). The previous works [27, 29] treat the MPM framework as a pre-training strategy, followed by fine-tuning on specific downstream tasks, such as classification or segmentation. In contrast, our PIC-G is an end-to-end model and does not adhere to the “pre-training then fine-tuning” training paradigm. Therefore, if we continue to use the center points masked out from the target point clouds as the input for positional embedding while using the MPM framework, we will not be able to provide the coordinates of the center points of the target point clouds for this positional embedding during inference, or it will lead to information leakage.
@@
Joint Sampling (JS) Module. Unlike pixels in 2D images that are aligned on the grid, 3D points in the point cloud are scattered and unordered [38], as illustrated in Fig 2 (b). Therefore, direct sampling or grouping based on location will lead to inconsistent results over input and target. The operations can become meaningful only when the correspondence between input and target point clouds can be established. To achieve this, they need to operate on the same set of point indices across inputs and targets instead of applying them to the same spatial region. To handle the above issues, we collect $N_{C}$ central points from each input point cloud and retrieve their indices, which we then use to obtain the center points of every patch in both the input and target point clouds. The process is shown in Fig. 2 (c). The key of our JS module is the consistency between center point indices of corresponding patches in both target and input point clouds. In other words, the order of the input token sequence and the target token sequence are well-aligned after being processed by the JS module. Such a design compensates for the missing positional embedding of the target while avoiding information leakage. Therefore, it facilitates the model in learning the inherent association between inputs and targets and streamlines the learning process. Subsequently, all point clouds search for neighborhoods containing $M$ points based on the center points corresponding to each patch.
@@
Model Architecture. We use a standard transformer with an encoder-decoder structure as the backbone of our Point-In-Context, and a simple $1\times 1$ convolutional layer as the task head for point cloud reconstruction. Inspired by Painter [13] and MAE [26], we explore two different baselines for PIC-G, naming them PIC-G-Sep and PIC-G-Cat. For PIC-G-Sep, we take the input and masked target point clouds parallel to the transformer and then merge their features after several blocks, using a simple average operation. For PIC-G-Cat, we concatenate the input and target to form a new point cloud. Then, we mask it globally and feed it to the transformer for prediction. We denote the prompt pair as $R^{k}_{i}=\{P_{i},T^{k}_{i}\}$ and query inputs, $Q^{k}_{j}=\{P_{j},T^{k}_{j}\}$, then PIC-G-Sep and PIC-G-Cat can be formalized as:
$P^{\text{Sep}}=\text{Transformer}([P_{i}\parallel P_{j}],([T^{k}_{i}\parallel T^{k}_{j}],M)),$ (1)
$P^{\text{Cat}}=\text{Transformer}([P_{i}\parallel T^{k}_{i}\parallel P_{j}\parallel T^{k}_{j}],M),$ (2)
@@
Loss Function. The model is trained to reconstruct the masked point patches. To this end, we use the $\ell_{2}$ Chamfer Distance [104] (CD) as the training loss. Specifically, we calculate the Chamfer Distance between each predicted patch $P$ and its corresponding ground truth $G$, which can be formulated as:
$\mathcal{L}_{CD}(P,G)=\frac{1}{\left|P\right|}\sum_{p\in P}\min_{g\in G}\|p-g\|_{2}^{2}+\frac{1}{\left|G\right|}\sum_{g\in G}\min_{p\in P}\|p-g\|_{2}^{2}.$ (3)
4.2 Point-In-Context-Segmenter Model

### v2

@@ 3.1 Modeling In-Context Learning in 3D Point Cloud
In-Context Learning in 2D. Previous works, such as Visual Prompt [visualprompt] and Painter [painter] combine two pairs of images that perform the same task into a grid-like image and randomly mask portions, following MAE [mae]. During training, these models [visualprompt, painter] take two pairs of images $I_{i}$ and $I_{j}$ as inputs and form two pairs, including a reference pair (also termed prompt), $R^{k}_{i}=\{I_{i},T^{k}_{i}\}$ and a query pair, $Q^{k}_{j}=\{I_{j},T^{k}_{j}\}$. $R^{k}_{i}$ is the task prompt containing one image $I_{i}$ and its corresponding target $T^{k}_{i}$. $Q^{k}_{j}$ represent the current input image $I_{j}$ and its target $T^{k}_{j}$, which perform the same task with the prompt pair. Here, $k$ represents the task index. When $i=j$, we term the prompt as an ideal prompt. During inference, only task example pairs and a query image are provided. They are combined into a grid-like image with a quarter mask to mask $T^{k}_{j}$, and a pre-trained model is used to restore the missing parts, as shown in Fig. 1 (b).
Modeling In-Context Learning in 3D. Following the same spirit of 2D in-context learning, for the first time, we design a similar procedure for 3D in-context learning. During training, each input sample contains two pairs of point clouds, including prompt point clouds $R^{k}_{i}=\{P_{i},T^{k}_{i}\}$ and query point clouds $Q^{k}_{j}=\{P_{j},T^{k}_{j}\}$, which perform the same task as in 2D in-context learning. Each pair consists of an input point cloud and its corresponding target point cloud for the given task. Similar to PointMAE [pointmae], we adopt the farthest point sampling (FPS) and K-nearest neighbor (KNN) techniques to convert the point clouds into a sentence-like data format. These point patches are subsequently encoded into tokens. During inference, the input point cloud is a combination of prompt and query point cloud, while the target point cloud consists of an example target along with masked tokens, as shown in Fig. 1(c) and Fig. 1(d). Consequently, we obtain predictions corresponding to the position of the mask. Given an input, the models can output corresponding targets following various prompts.
@@ 3.2 ShapeNet In-Context Datasets
To establish the first benchmark for 3D in-context learning, we carefully curate datasets and define task specifications, as there is no prior benchmark. Initially, we collect samples from representative datasets like ShapeNet [shapenet] and ShapeNetPart [shapenetpart], transforming them into the “input-target” format as stated in Sec. 3.1. To augment the sample size for part segmentation, we apply random operations such as point cloud perturbation, rotation, and scaling to ShapeNetPart. This yields an extensive dataset with 217,454 samples across four tasks. Each sample includes an input point cloud and its corresponding task-specific target. We standardize the XYZ coordinates of inputs and outputs to the range between [1,1][-1,1], and unify all samples as pure point clouds in $(N,3)$ shape.
@@
Registration. This task aims to restore a rotated point cloud to its initial orientation. We assume that during both training and inference, the query point cloud and the prompt are synchronized in terms of the rotation angle. To prevent interference with the denoising and reconstruction tasks, we configure the registration task’s output to include both an upright and an upside-down point cloud. We also introduce five corruption levels for the evaluation on this task.
@@
3.3 Multi-Entity In-Context Datasets
@@
To further improve the performance and generalization of PIC to segmentation tasks, we collect more datasets related to part segmentation in point clouds. These datasets include ShapeNetPart [shapenetpart], originally used for synthetic object part segmentation, Human3D [human3d] for human body-part segmentation, BEHAVE [behave] for human-object interaction segmentation, and AKB-48 [akb48] for real-collected object part segmentation. These datasets contain point clouds widely used in 3D human-machine interaction scenarios [nam2023cyclic], including human, object, and human-object interaction point clouds. We process these datasets into a format suitable for part segmentation and integrate them into large-scale Multi-Entity In-Context Datasets, consisting of approximately 140k samples. Moreover, to test the model’s generalization, we exclude AKB-48 from the training set and use it as a one-shot generalization test dataset.
@@
ShapeNetPart [shapenetpart] dataset consists of 16,881 shapes categorized into 16 classes, with 50 part labels. Within each class, the number of parts varies from 2 to 6. As described in Sec. 3.2, we aim to augment the challenge of the ShapeNetPart dataset and enhance its representation of real-world point cloud scenarios. To achieve this, we apply various random operations to the dataset, introducing noise to simulate imperfect point clouds commonly encountered in real-world environments. Then, we transform all part segmentation labels of a sample into points in space, each point containing XYZ coordinates. These transformed points are then aggregated to construct a point cloud with the same shape as the input. Finally, we combine the input with the labeled point cloud to create a sample in our dataset.
Figure 2: (a) The pre-training pipeline used in previous works. When performing Masked Point Modeling (MPM), these works [pointbert, pointmae, pointm2ae] use the center position of the target patches for position embedding, which results in information leakage. (b) Difference between 2D images and 3D point clouds. For 2D images, the semantic information at corresponding pixel positions in the input and target images is consistent. However, after performing grouping operations on 3D point clouds, the order of points will inevitably be disrupted, resulting in inconsistency between the point sequences of the input and output point clouds. (c) Joint Sampling module involves recording the indices of sampled center points and employing the K-nearest neighbor strategy to sample both the input and target point clouds concurrently.
@@
Human3D [human3d] is a dataset comprising point cloud videos captured in real-world scenarios, annotated based on human body parts. Each sample of it contains two persons, both divided into 15 parts. To seamlessly integrate prompts into our proposed in-context learning framework, we divide each sample containing two individuals into two separate samples. Each sample focuses on one person with annotations for human-part segmentation. Given the complexity and difficulty in distinguishing between the original 15 body parts, we merge them into 6 categories: head, body, left arm, right arm, left leg, and right leg. This consolidation aims to enhance the semantic representation of human body parts. Additionally, we treat each frame of the dynamic video as a static sample in our dataset. This approach allows us to create a single-person human body part segmentation dataset, comprising 57,078 samples.
@@
BEHAVE [behave] is an extensive dataset focusing on full-body human-object interactions, featuring multi-view RGBD frames paired with corresponding 3D SMPL and object fits. Additionally, it includes annotated contact points between humans and objects. The BEHAVE dataset encompasses roughly 15K frames capturing 8 subjects engaging in various interactions with 20 commonly encountered objects. Presented in video format, each frame of the RGBD video is transformed into point clouds, and treated as an individual sample within our dataset. These point cloud samples are annotated for segmentation, covering both the person and the object. After gathering around 15k point cloud samples, we excluded samples in which the proportion of points attributed to objects is smaller than 5% of the entire point cloud sample.
@@
AKB-48 [akb48] is a comprehensive-articulated object dataset containing 2,037 real-world 3D models across 48 categories. Each model is meticulously scanned from its real-life counterpart and further refined manually. Throughout the data collection, AKB-48 categorizes objects into multiple components, facilitating their utilization in part segmentation tasks. However, some categories in AKB-48 inherently pose challenges due to semantic ambiguity or confusion. For example, within the “toy” category, samples may include both dinosaurs and cars, which do not share identical labels, highlighting semantic ambiguity. Another instance is the “folding rack” category, which presents a symmetric object segmented into two or even four components despite its identical left and right sides, showcasing semantic confusion. To address these issues, we conduct filtering and preprocessing operations on the AKB-48 dataset. Firstly, categories prone to semantic confusion are removed. Furthermore, we manually consolidate components with similar semantics into a unified category. Finally, following a similar approach to the ShapeNetPart dataset, each component is labeled and merged to form a comprehensive point cloud with segmentation labels. Given AKB-48’s diverse component categories and absence of overlap with previous datasets, we consider it a one-shot segmentation generalization test dataset.
@@ 4 Methodology
In this section, we first introduce our Point-In-Context (PIC), a general point cloud framework capable of handling multiple tasks via the in-context learning capability. Next, to enhance the performance and generalization of PIC in segmentation tasks, we highlight the extended version of PIC, termed PIC++, which not only retains the multitask capabilities but also excels in part segmentation across multiple datasets.
@@
Figure 3: Overall scheme of our Point-In-Context (PIC). Top: Training pipeline of the Masked Point Modeling (MPM) framework. During training, each sample comprises two pairs of input and target point clouds that tackle the same task. These pairs are fed into the transformer model to perform the masked point reconstruction task, which follows a random masking process. Bottom: In-context inference on multitask. Our PIC could infer results on various downstream point cloud tasks, including reconstruction, denoising, registration, and part segmentation.
4.1 Point-In-Context Model
@@
MPM for 3D In-Context Learning. Building upon previous works [visualprompt, painter], we adopt Masked Point Modeling (MPM) for the training of point cloud understanding and introduce Point-In-Context (PIC). In PIC, we conceptualize the local area of point clouds within the MPM as akin to image tokens within the MIM, thereby enabling us to harness the power of the transformer architecture for processing both point cloud and image data seamlessly. Illustrated in Fig. 3, our training process aims to reconstruct the masked point patches, facilitated by the presence of input point clouds and visible target point cloud patches. During the inference stage, given a query point cloud, PIC will perform the specific task: reconstruction, denoising, registration, or part segmentation, according to the prompt. Subsequently, it generates target outputs corresponding to the query according to the task demonstrated by the prompt.
@@
Information Leakage. The conventional 3D MPM approach of employing position embeddings may introduce information leakage. That is because previous pre-training pipelines embed positional information with the sampled center point coordinates of all patches, even for those that are masked out (invisible), as shown in Fig. 2 (a). The previous works [pointmae, pointm2ae] treat the MPM framework as a pre-training strategy, followed by fine-tuning on specific downstream tasks, such as classification or segmentation. In contrast, our PIC is an end-to-end model and does not adhere to the “pre-training then fine-tuning” training paradigm. Therefore, if we continue to use the center points masked out from the target point clouds as the input for positional embedding while using the MPM framework, we will not be able to provide the coordinates of the center points of the target point clouds for this positional embedding during inference, or it will lead to information leakage.
Figure 4: (a) Comparison of generating targets between In-Context Labeling (PIC++) and pre-defined label map (PIC). $P$ represents the number of all parts ($P\gg N_{B}>|C_{i}|$). (b) Illustration of In-Context Enhancing. We first randomly apply common corruptions on a clean point cloud and construct in-context pairs. Then we train our PIC++ on the combined training set of Multi-Entity In-Context Datasets and the constructed in-context pairs.
@@
Joint Sampling (JS) Module. Unlike pixels in 2D images that are aligned on the grid, 3D points in the point cloud are scattered and unordered [pointnet], as illustrated in Fig 2 (b). Therefore, direct sampling or grouping based on location will lead to inconsistent results over input and target. The operations can become meaningful only when the correspondence between input and target point clouds can be established. To achieve this, they must operate on the same set of point indices across inputs and targets instead of applying them to the same spatial region. To handle the above issues, we collect $N_{C}$ central points from each input point cloud and retrieve their indices, which we then use to obtain the center points of every patch in both the input and target point clouds. The process is shown in Fig. 2 (c). The key of our JS module is the consistency between center point indices of corresponding patches in both target and input point clouds. In other words, the order of the input token sequence and the target token sequence are well-aligned after being processed by the JS module. Such a design compensates for the missing positional embedding of the target while avoiding information leakage. Therefore, it facilitates the model in learning the inherent association between inputs and targets and streamlines the learning process. Subsequently, all point clouds search for neighborhoods containing $M$ points based on the center points corresponding to each patch.
@@
Model Architecture. We use a standard transformer with an encoder-decoder structure as the backbone of our Point-In-Context, and a simple $1\times 1$ convolutional layer as the task head for point cloud reconstruction. Inspired by Painter [painter] and MAE [mae], we explore two different baselines for PIC, naming them PIC-Sep and PIC-Cat. For PIC-Sep, we take the input and masked target point clouds parallel to the transformer and then merge their features after several blocks, using a simple average operation. For PIC-Cat, we concatenate the input and target to form a new point cloud. Then, we mask it globally and feed it to the transformer for prediction. We denote the prompt pair as $R^{k}_{i}=\{P_{i},T^{k}_{i}\}$ and query inputs, $Q^{k}_{j}=\{P_{j},T^{k}_{j}\}$, then PIC-Sep and PIC-Cat can be formalized as:
$P^{\text{Sep}}=\text{Transformer}([P_{i}\parallel P_{j}],([T^{k}_{i}\parallel T^{k}_{j}],M)),$
$P^{\text{Cat}}=\text{Transformer}([P_{i}\parallel T^{k}_{i}\parallel P_{j}\parallel T^{k}_{j}],M),$
@@
Loss Function. The model is trained to reconstruct the masked point patches. To this end, we use the $\ell_{2}$ Chamfer Distance [chamferdistance] (CD) as the training loss. Specifically, we calculate the Chamfer Distance between each predicted patch $P$ and its corresponding ground truth $G$, which can be formulated as:
$\begin{split}\mathcal{L}_{CD}(P,G)=\frac{1}{\left|P\right|}\sum_{p\in P}\min_{g\in G}\|p-g\|_{2}^{2}+\\ \frac{1}{\left|G\right|}\sum_{g\in G}\min_{p\in P}\|p-g\|_{2}^{2}.\end{split}$ (3)
4.2 Point-In-Context++ Model